the molar solubility of cobalt (II) carbonate in a water soution is ???
1
Expert's answer
2018-03-23T08:47:55-0400
The Ksp of cobalt(II) carbonate, CoCO3, is 1.0·10-10.

CoCO3 ↔ Co2+ + CO32-

1 mole ↔ 1 mole + 1 mole

x moles ↔ x moles + x moles

Ksp = [Co2+]·[CO32-]

Ksp = 1.0·10-10 = (x)*(x) = x2

x = (1.0·10-10)1/2

x = 1·10-5 M

The molar solubility of cobalt (II) carbonate in a water solution is 1·10-5 M.